A business uses two 3kW fires an average of 20 hours each per week and six 150W lights for 30hours each per week. If the cost of electricity is 14.25p per unit, determine the weekly cost of electricity to the business.
1
Expert's answer
2020-05-04T12:49:50-0400

First of all calculate the total no. of units

total units ="3\times20+0.15\times6\times30=60+27=87"


total cost = no. of units * 14.25 = 87*14.25 = 1239.75